OpenStudy (anonymous):

which set of data has a mean of 6 and mode of 5 ?

OpenStudy (anonymous):

This is one: {5, 5, 5, 8, 7}

OpenStudy (anonymous):

Here another one {5, 5, 5, 5, 8, 8}

OpenStudy (anonymous):

Here is another one {5, 5, 5, 5, 5, 8, 8, 7} You can construct infinitely many.

OpenStudy (anonymous):

option are: A. 5, 12 , 1, 5, 7 B. 3, 8, 5, 9, 3 C.2, 11, 5, 9, 3 D. 10, 7, 5, 8, 5.

OpenStudy (anonymous):

i don't know which is ?

OpenStudy (anonymous):

which of the above datas have mode 5?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

\[ \frac{5 + 12 + 1 + 5 + 7}5=\frac {30}5 =6 \]

OpenStudy (anonymous):

It is clear that the mode is 5. Since 5 is repeated twice

OpenStudy (anonymous):

and the others are repeated once.

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

yes. thank you so much. professor !!

OpenStudy (anonymous):

mean is the average of all the numbers ie add all the numbers and divide them by their quantity. Mode is the the number that repeats the most in the series . If no number repeats there is no mode. So in these options you can start by looking for series with mode = 5 as kit will be easy to find it. It comes as 5 for A and D. if you calculate median for A it is \[(5+12+1+5+7) \div 5 = 305 \div 5 = 6\] and this indeed is the answer.

OpenStudy (anonymous):

it is \[30 \div 5 = 6\]

OpenStudy (anonymous):

thank you Panchtatvan !!!

OpenStudy (anonymous):

yes 30/5=6